mathslover (mathslover):

this was asked earlier but didn't get the answer \[\large{\frac{\sqrt{\sqrt{5}+2}+\sqrt{\sqrt{5}-2}}{\sqrt{\sqrt{5}+1}}}\]

mathslover (mathslover):

wait editing the latex

OpenStudy (anonymous):

done ask the question reapeted http://openstudy.com/code-of-conduct

mathslover (mathslover):

@best.shakir it is not against code of conduct: Posting a question multiple times within a short period of time. i had not posted withing a short period of time

OpenStudy (anonymous):

i m sorry

mathslover (mathslover):

\[\large{\frac{\sqrt{\sqrt{5}+2}+\sqrt{\sqrt{5}-2}}{\sqrt{\sqrt{5}+1}}}\]

mathslover (mathslover):

\[\large{\frac{\sqrt{\sqrt{5}+2}+\sqrt{\sqrt{5}-2}}{\sqrt{\sqrt{5}+1}}-\sqrt{3-2\sqrt{2}}}\]

OpenStudy (anonymous):

it is 1

OpenStudy (anonymous):

 u might have done a mistake here